OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = 1/2 and b = 3.

OpenStudy (midhun.madhu1987):

y = mx + b where m = 1/2 and b = 3

OpenStudy (anonymous):

no standard form is ax+by=c @midhun.madhu1987

OpenStudy (midhun.madhu1987):

You can still rearrange y=mx + c to ax+by+c=0

OpenStudy (anonymous):

y = (1/2)x + 3 (subtract (1/2)x from both sides) (-1/2)x + y = 3 (multiply both sides by -2) i think its x - 2y = -6

OpenStudy (anonymous):

@dstinefield thanks!

OpenStudy (anonymous):

no problem